Write the equations for the following functions in vertex form:
y=x^2−14x+30
y=x^​2+7x+10

Answer by lwsshak3(11628)

Write the equations for the following functions in vertex form:
y=x^2−14x+30
complete the square:
y=(x^2-14x+49)-49+30
y=(x-7)^2-19
..
y=x^​2+7x+10
y=(x^​2+7x+49/4)-49/4+10
y=(x+7/2)^2-49/4+40/4
y=(x+7/2)^2-9/4